OpenStudy (anonymous):

On a new temperature scale (°Y), water boils at 155.0 °Y and freezes at 0.00 °Y. Calculate the normal human body temperature using this temperature scale. On the Fahrenheit scale, normal human body temperature is 98.6 °F, and water boils at 212.0 °F and freezes at 32.0 °F. How do you start this and what kind of formulas do you need to solve it? I need a breakdown please

OpenStudy (aaronq):

build a ratio. For the first part: \(\dfrac{x}{155^oY}=\dfrac{98.6 °F}{BP\;of\;water\;in\;F}\) where x is the "normal human body temperature" in Y degrees.

OpenStudy (aaronq):

okay wait i didn't know the freezing point of water was 32 F, that changes things. hold on

OpenStudy (aaronq):

so: \(\dfrac{x}{155^oY-0^oY}=\dfrac{98.6 °F}{212^oF-32^oF}\) what you wanna do is basically scale it but building a ratio with the end points

OpenStudy (anonymous):

That makes sense!

OpenStudy (aaronq):

honestly, no one in science uses Fahrenheit.

OpenStudy (anonymous):

true. now if we had to do this in celsius it would be in the same way?

OpenStudy (anonymous):

except just different numbers since freezing is 0 in celsius and boiling is 100

OpenStudy (anonymous):

now i got x = 84.91 why does that not look correct, or do i have to convert it?

OpenStudy (aaronq):

for celsius it would be the same.

OpenStudy (anonymous):

i have a website with the answer to this question (just no work shown) and the answer they got is 57.3

OpenStudy (aaronq):

i got x=84.90555 and it makes sense, because the F scale is broken up into 180 units the Y scale is broken into 155 units so each Y unit (degree) is accounts a larger temperature range than the F degree

OpenStudy (anonymous):

what exactly do you mean by 180 units, wouldn't it be 212 units?

OpenStudy (aaronq):

in terms of water, no. from the freezing point (32) to the boiling point of water (212) there are (21-32) 180 units.

OpenStudy (aaronq):

the fictitious Y scale (just like the celsius scale) take the freezing point of water as zero

OpenStudy (anonymous):

Ohh i get it. Hmm so that 57.3 answer is wrong?

OpenStudy (aaronq):

yes

OpenStudy (anonymous):

you see the attachment? seems reliable since it's from the textbook

OpenStudy (aaronq):

i think they messed up. if you do the same exercise with the celsius scale: \(\dfrac{x}{155-0}=\dfrac{37}{100-0}\) x=57.35

OpenStudy (anonymous):

Oh wow didn't even think about that. Interesting. thank you very much!

OpenStudy (aaronq):

you'd be surprised how many mistakes there are in books (specially in the answers section). no problem !

OpenStudy (anonymous):

so would this way of doing it for these sorts of questions always work?

OpenStudy (aaronq):

yeah. building ratios is useful in many types of questions, not only this type. It works when you're comparing to things and they're on fixed scales.

OpenStudy (aaronq):

two things*

OpenStudy (anonymous):

hmm how about for this one then, because i just tried doing it and it's off from the answer given (even though i'm skeptical since there are errors)

OpenStudy (anonymous):

i get 29.6

OpenStudy (anonymous):

oh wait i actually get it. they added the 40 degrees since it takes 40 degrees to make the scales equal to each other

OpenStudy (aaronq):

yeah, i was gonna say, it starts at 40.

OpenStudy (anonymous):

and because in the previous problem it started at 0 we didn't have to add or subtract right?

OpenStudy (aaronq):

I just thought about that, and it doesn't seem like it. 85-32=53, which is far from the answer.

OpenStudy (anonymous):

true. alright then. i'll just have to think more critically with these questions! haha thank you once again

OpenStudy (aaronq):

damn, i'm sorry. you did have to. look: you'd have to subtract 32F scaled to the Y temperature x/155=32/180 -> x=27.55555 85-27.55555=57.444444

OpenStudy (anonymous):

where did you get the 85 from?

OpenStudy (aaronq):

from the initial calculation it was 84.91, i rounded for simplicity.

OpenStudy (anonymous):

So you have to perform 2 calculations using the same equation and then subtracting the both answers to get the real one?

OpenStudy (aaronq):

yeah, it seems like it.

OpenStudy (anonymous):

alright, thank you

OpenStudy (aaronq):

no problem. sorry about the errors earlier.

OpenStudy (anonymous):

not a problem!